OpenStudy (anonymous):

Really need help with exponential functions! PLEASE 1/25 = 5^x+4 8 = 2^x+4 (sqrt6)^8x = 216^x-3

OpenStudy (anonymous):

8=2^x +4 move the 4 and subtract it to 8 so is 4=2^x and then do logarithm

OpenStudy (anonymous):

so the answer for the first one is 2?

OpenStudy (anonymous):

or second one***

OpenStudy (anonymous):

I was doing the second one

OpenStudy (anonymous):

the first one is the same

OpenStudy (agent0smith):

4=2^x you don't need to take a log for this. 2 to what power equals 4... what's 2^2?

OpenStudy (anonymous):

true ^

OpenStudy (agent0smith):

1/25 = 5^x+4 subtract 4 from both sides first

OpenStudy (anonymous):

i just don't understand how to do it because of the exponent

OpenStudy (anonymous):

where there is a power to the exponents, you use logarithms

OpenStudy (agent0smith):

Don't worry about the exponent, until you've got it alone on one side... then deal with it.

OpenStudy (anonymous):

i'm sorry i still don't get it :(

OpenStudy (agent0smith):

Just worry about getting it alone for now. Subtract 4 on both sides of this: 1/25 = 5^x+4

OpenStudy (anonymous):

-3 24/25 = 5^x

OpenStudy (agent0smith):

It'd be easier to make it into a decimal -3.96 = 5^x this has no solution - an exponential like 5^x is always positive, your answer is no solution.

OpenStudy (anonymous):

thats not even a possible answer :(

OpenStudy (anonymous):

the possible answers are x= -7/2 x = -6 x = 9/2 x = 2

OpenStudy (agent0smith):

are you sure this is the equation? 1/25 = 5^x+4

OpenStudy (anonymous):

the x+4 is all ^

OpenStudy (anonymous):

like x+4 is all an exponent

OpenStudy (agent0smith):

You gotta specify that, use brackets... 1/25 = 5^(x+4) 1/25 is the same as 5^(-2) \[\Large 5^{-2}=5^{x+4}\]

OpenStudy (anonymous):

oh okay sorry i will next time

OpenStudy (anonymous):

so 8 = 2^(x+4) and (sqrt 6)^8x = 216^(x-3)

OpenStudy (agent0smith):

Wait we didn't finish the last one... did you find x?

OpenStudy (anonymous):

no i don't understand how to do it from exponents at all

OpenStudy (agent0smith):

\[\Large 5^{-2}=5^{x+4}\]since the bases, 5, are equal, the exponents must be equal -2 = x+4

OpenStudy (anonymous):

so -6

OpenStudy (anonymous):

how do i do the other ones?

OpenStudy (agent0smith):

(sqrt 6)^8x = 216^(x-3) sqrt 6 is 6^0.5, 216 is 6^3 so we have \[\Large (6^{0.5})^{8x} = (6^3)^{x-3}\]

OpenStudy (agent0smith):

you can simplify that if you remember some exponent rules